OpenStudy (anonymous):

How do you solve these? Solve for t. l m F= ---- t^2

OpenStudy (anonymous):

\[F = \frac{lm}{t^2}\]\[Ft^2 = lm\]\[t^2 = \frac{lm}{F}\]\[t = \sqrt{\frac{lm}{F}}\]